RON Ltd has the following capital structure components:

  • Five million shares issued with a current market price of 11. Equity holders require a 8% return.
  • $10 million face value of Corporate bonds outstanding. These bonds pay an annual coupon of 6% and currently trade at a yield to maturity of 6%.

If the firm faces a corporate tax rate of 30%, compute RON Ltd's Weighted Average Cost of Capital (WACC). Enter your answer in decimal form to FOUR decimal places. For example 10.34%, would be entered as 0.1034
Answer #1

Number of Common Shares = N = 5 million and Price per Share = P = $ 11, Required Return = 8%

Market Value of Common Stock = 5 x 11 = $ 55 million

Face Value of Debt = $ 10 million, Bond Coupon = 6 % and Yield = 6%, As bond coupon equals the yield, the market value of debt should equal its face value.

Hence, Debt Proportion = 10 / (10+55) = 0.15385 and Equity Proportion = 55/65 = 0.84615

Tax Rate = 30 %

Cost of Debt = Yield = 6%

Therefore, WACC = 0.84615 x 8 + (1-0.3) x 0.15385 x 6 = 7.41538 % or 0.0741